TOWARD A COMPLETE STRUCTURE OF BUTADIENE WITH HIGH-RESOLUTION INFRARED SPEC-TROSCOPY

\underline{NORMAN C. CRAIG}, JEDIDIAH L. DAVIS, and KEVIN J. WEIDENBAUM, Department of Chemistry, Oberlin College, Oberlin, OH 44074; MICHAEL LOCK, Physikalisch-Chemisches Institut, Heinrich-Buff-Ring 58, Justus-Liebig-Universität, D-35392 Giessen, Germany.

Although the structure of butadiene is used in initial discussions of delocalized π-bonding in every organic textbook, the structure of this fundamental molecule is incompletely known. Information about the CCCC backbone comes from electron diffraction studies.^{1} Caminati et al have proposed a partial structure from this data, a low-resolution infrared investigation of butadiene,^{2} and their MW investigation of butadiene-1,1-d_{2}.^{3} We have begun a high-resolution (0.002 cm^{−1}) infrared investigation of the complete structure of this nonpolar molecule, which is MW-silent. For the normal species, a C-type band due to out-of-plane CH flapping at 908.072 cm^{−1} and a perturbed A-type band due to antisymmetric C=C stretching at 1596.446 cm^{−1} have been analyzed. The ground state rotational constants are A = 1.3903839(10), B = 0.1478862(2), and C = 0.1336942(2) cm^{−1} for a Watson-type Hamiltonian. For the 2,3-d_{2} species, the C-type band at 908.044 cm^{−1} has been analyzed to give A = 1.0203626(5), B = 0.1472767(3), and C = 0.1287171(3) cm^{−1} for the ground state rotational constants. Structural implications of these data will be discussed. Currently, we are attempting a stereospecific synthesis of the two 1,4-d_{2} isotopomers and have hopes of obtaining the two singly-substituted 13C isotopomers.

(1) Kveseth, K.; Seip, R.; Kohl, D. A. Acta Chem. Scand. A 1980, 34, 31.

(2) Cole, A. R. H.; Mohay, G. M.; Osborne, G. A. Spectrochim. Acta 1967, 23A, 909.

(3) Caminati, W.; Grassi, G; Bauder, A. Chem. Phys. Letters 1988, 148, 13.